OpenStudy (anonymous):

Given that f(x)= absolute(2x+6) determine where it is not differentiable?

OpenStudy (anonymous):

at the corner

OpenStudy (anonymous):

not diff at 0

OpenStudy (anonymous):

oh yes it is

OpenStudy (anonymous):

nah man it isnt

OpenStudy (anonymous):

sure it is

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

absolute value of 2x + 6 =2x+6 if 2x+6 > 0 i.e if x > -3

OpenStudy (anonymous):

and of course it is -2x-6 if x < -3

OpenStudy (anonymous):

is continuous at -3, but has a corner there, so not differentiable at -3

OpenStudy (anonymous):

-2x-6 if x < -3

OpenStudy (anonymous):

yah man

OpenStudy (anonymous):

@akileez here is a nice picture of the corner http://www.wolframalpha.com/input/?i= |2x%2B6|

OpenStudy (anonymous):

i am confused on this one now?

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

\[|2x+6|\] is actually a piece-wise defined function, just as \[|x|\] is

OpenStudy (anonymous):

ok gotcha

OpenStudy (anonymous):

if what is inside the absolute values sign it positive, then it is just itself. for example if x = 2 then\[ |2\times 2 +6|=2\times 2+6\]

OpenStudy (anonymous):

on the other hand if what is in the absolute value sign in negative, it is the opposite of that

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I understand absolute vlaue

OpenStudy (anonymous):

it's the left and right calc that gets me

OpenStudy (anonymous):

for example if x = -5 then \[|2\times -5+6|=-2\times- 5 - 6\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

in other words \[|2x+6|=2x+6\] or \[|2x+6|=-2x-6\]

OpenStudy (anonymous):

so you have two different lines, one with slope 2 and the other with slope -2. these are different, and they are your derivatives

OpenStudy (anonymous):

finally! that is the right and left side calc correct?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

left being the minus 2x

OpenStudy (anonymous):

and where do they change?

OpenStudy (anonymous):

great I have it meow!

OpenStudy (anonymous):

they change when \[2x+6>0 \] \[2x>-6\] \[x>-3\]

OpenStudy (anonymous):

in other words they change at -3, so this function is not differentiable there

OpenStudy (anonymous):

dose that mean at -3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

did you look at the picture i sent? you see a corner at -3

OpenStudy (anonymous):

why in the left side calc is the 6 a minus? I thought that the only x would change signs?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

it only opened to the home page with no equation, pic or anything else

OpenStudy (anonymous):

|x| is x if x >0 and -x if x < 0 ok with that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this means that |2x+6| = 2x+6 if 2x+6 > 0 or -2x-6 if 2x+6<0

OpenStudy (anonymous):

i just replaced the "x" by "2x+6"

OpenStudy (anonymous):

ok with that?

OpenStudy (anonymous):

yes i see, so both signs must change - makes sense! thanks!

OpenStudy (anonymous):

but of course you have to answer in terms of x, not in terms of 2x+6

OpenStudy (anonymous):

so if 2x+6>0 that means x >-3 and that is where this changes from 2x+6 to -2x-6

OpenStudy (anonymous):

that is where the slope changes from -2 to 2

OpenStudy (anonymous):

and so that is where your function is not differentiable.

OpenStudy (anonymous):

gotcha!